3. (10 points) Based on a random sample of size n from the pdf below, derive the MLE of o. , 00
Answer #1

from above L(\theta)=f(x1)*f(x2)*f(x3)*...*f(xn)=(\theta+1)n*(x1*x2*..xn)\theta

Ln(L(\theta))=n*Ln(\theta+1)+\theta*\sumln(xi)

(d/d\theta)*Ln(L(\theta)) =n/(1+\theta)+\sumln(xi)

putting  (d/d\theta)*Ln(L(\theta)) equal to 0;

n/(1+\theta) =-\sumln(xi)

\theta =-1-n/(\sumln(xi))